# Recursive variational Gaussian approximation with the Whittle likelihood for linear non-Gaussian state space models

### v2

@@ 3.4 Real univariate and bivariate examples
We apply R-VGA-Whittle to estimate parameters in univariate and bivariate stochastic volatility models from real data. We first apply it to daily JPY-EUR exchange rate data (Kastner et al.,, 2017) from April 1st, 2005 to August 6th, 2015, which consists of $2650$ observations. Typically, the data used in SV models are de-meaned log-returns, which are computed from the raw exchange rate values as:
@@
where $r_{t}$ is the exchange rate on day $t$, and $y_{t}^{*}$ is the log-return on day $t$. We fit the model (23)–(26) to this time series of de-meaned log-returns with $T=2649$, and use the initial variational distribution specified in (31).
The HMC-exact trace plots in Figures 11(a) of the online supplement show that posterior samples of $\phi$ are relatively well-mixed, but that the posterior samples of $\sigma_{\eta}$ are highly correlated. We find it necessary to thin these samples by a factor of 100 in order for them to be considered uncorrelated. In contrast, HMC-Whittle produces well-mixed posterior samples, so that no thinning is needed.
We compare the posterior densities from R-VGA-Whittle, HMC-Whittle and HMC-exact in Figure 9. The posterior densities obtained from R-VGA-Whittle are similar, but slightly wider than those obtained from HMC-Whittle and HMC-exact.
Figure 9: Posterior distributions from HMC-exact (blue), and approximate posterior distributions from R-VGAL-Whittle (red) and HMC-Whittle (yellow) for the univariate SV model fitted using JPY-EUR exchange rate data with $T=2649$. Diagonal panels: Marginal posterior distributions of the parameters. Off-diagonal panels: Bivariate posterior distributions of the parameters.
@@
Figures 12(a)12(c) in Section S5 of the online supplement show the HMC-exact and HMC-Whittle trace plots for the bivariate SV model with exchange rate data. Similar to the simulation study, we find that HMC-exact samples are highly correlated; samples of parameters in $\mathbf{\Phi}$ tend to be less correlated than samples of parameters in $\mathbf{\Sigma}_{\eta}$. We thin the HMC-exact samples by a factor of 50 so that the remaining samples can be treated as uncorrelated. Posterior samples from HMC-Whittle have low correlation and are mixed well for all parameters, so no thinning is necessary.
@@
Table 1 shows that R-VGA-Whittle is twice as fast as HMC-Whittle in the univariate case, and more than 80 times faster than HMC-Whittle in the bivariate case. R-VGA-Whittle is significantly faster than HMC-exact, with approximately a 260-fold speedup in the univariate case and a 700-fold speedup in the bivariate cases.
Figure 10: Posterior distributions from HMC-exact (blue), and approximate posterior distributions from R-VGAL-Whittle (red) and HMC-Whittle (yellow) for the bivariate SV model fitted using GBP-EUR and USD-EUR exchange rate data with $T=2649$. Diagonal panels: Marginal posterior distributions of the parameters. Off-diagonal panels: Bivariate posterior distributions of the parameters.
@@ 4 Conclusion
We develop a novel sequential variational Bayes algorithm that uses the Whittle likelihood, called R-VGA-Whittle, for parameter inference in linear non-Gaussian state space models. The Whittle likelihood is a frequency-domain approximation of the exact likelihood. It is available in closed form and its log can be expressed as a sum of individual log-likelihood frequency components. At each iteration, R-VGA-Whittle sequentially updates the variational parameters by processing one frequency or a block of frequencies at a time. We also demonstrate the use of the Whittle likelihood in HMC (HMC-Whittle), and compare the parameter estimates from R-VGA-Whittle and HMC-Whittle to those of HMC with the exact likelihood (HMC-exact). We show that the posterior densities from the three methods are generally similar, though posterior densities from R-VGA-Whittle and HMC-Whittle tend to be wider compared to those from HMC-exact. As HMC-Whittle only targets the parameters and not the states, it tends to produce posterior samples of the parameters that are less correlated than those from HMC-exact. We also show that R-VGA-Whittle and HMC-Whittle are much more computationally efficient than HMC-exact. R-VGA-Whittle is particularly efficient when dealing with bivariate time series, where it can be up to two orders of magnitude faster than HMC-Whittle and three orders of magnitude faster than HMC-exact.
There are several avenues for future research. First, R-VGA-Whittle assumes a Gaussian variational distribution, so future work could attempt to extend this framework to more flexible variational distributions for approximating skewed or multimodal posterior distributions. Second, some tuning parameters of R-VGA-Whittle (such as the number of damping observations and number of damping steps) are currently user-defined, so an adaptive tuning scheme could be developed to select these parameters. Third, as the Whittle likelihood is known to produce biased parameter estimates for finite sample sizes, using R-VGA-Whittle with the debiased Whittle likelihood of Sykulski et al., (2019) may improve parameter estimates.
@@ Acknowledgements
This work was supported by ARC SRIEAS Grant SR200100005 Securing Antarctica’s Environmental Future. Andrew Zammit-Mangion’s research was also supported by ARC Discovery Early Career Research Award DE180100203.
@@
Recall that the periodogram is an estimate of the spectral density and can be viewed as a measure of the relative contributions of the frequencies to the total variance of the process (Percival and Walden,, 1993). For example, Figure S1 shows the periodograms for the linear Gaussian SSM in Section 3.1 and the univariate SV model in Section 3.2. This figure shows that in general, power decreases as frequency increases. Frequencies that carry less power tend to have less effect on the trajectory of the variational mean; see, for example, the trajectories of the parameters of the linear Gaussian model in Figure S2. The trajectory of $\phi$ moves towards the true parameter value within the first few frequencies, with only minor adjustments afterwards. The trajectories for $\sigma_{\eta}$ and $\sigma_{\epsilon}$ behave in a similar manner, but with more gradual changes. This behaviour can also be observed in Figure S3, in which the trajectories of both parameters in the univariate SV model exhibit large changes in the first few iterations, then stabilise afterwards. Higher frequencies that have little effect on the variational mean can thus be processed together in “blocks”. The R-VGA-Whittle updates can then be made based on the Whittle log-likelihood for a “block” of frequencies, which we compute as the sum of the individual log-likelihood contribution evaluated at each frequency within that block.
@@
Figure S1: Plot of raw periodograms of the data (black) and smoothed periodograms (light red) for two models. Vertical lines show the frequency cutoffs at half-power (3dB, red dashed line), one-fifth-power (7dB, blue dotted line) and one-tenth-power (10dB, purple dash-dotted line).
Figure S2: Trajectories of the variational mean for the parameters in the LGSS model estimated using damped R-VGA-Whittle (Algorithm 2) with no blocking. The true parameter values are marked with dashed lines.
Figure S3: Trajectories of the variational mean for the parameters in the univariate SV model estimated using damped R-VGA-Whittle (Algorithm 2) with no blocking. The true parameter values are marked with dashed lines.
It is important to choose an appropriate point to start blocking, as well as an appropriate length for each block of frequencies. To determine the frequency at which to begin blocking, we borrow a concept from physics and electrical engineering called the 3dB frequency cutoff, which is the frequency on the power spectrum at which the power drops to approximately half the maximum power of the spectrum. The value 3 comes from the fact that, on the decibel scale, a drop from a power level $P_{1}$ to another level $P_{2}=\frac{1}{2}P_{1}$ results in a drop of 3 decibels. The change in decibels is calculated as
@@
We employ a similar idea here, where we define the “cutoff” frequency as the frequency at which the power drops to half of the maximum power in a smoothed version of the periodogram. Specifically, we seek the frequency $\omega^{*}$ such that $\tilde{I}(\omega^{*})=1/2\max(\tilde{I}(\omega))$, where $\tilde{I}(\omega)$ is the periodogram smoothed using Welch’s method (Welch,, 1967). Briefly, Welch’s method works by dividing the frequencies into equal segments, computing the periodogram for each segment, and then averaging over these segments to obtain a smoother periodogram with lower variance. We also experiment with other cutoffs, such as the frequencies at which the power are one-fifth and one-tenth of the maximum power in the smoothed periodogram, which correspond to a drop of approximately 7dB and 10dB, respectively. We mark these cutoff frequencies in Figure S1.
@@
Figure S4: Posterior densities of the parameters of the linear Gaussian SSM estimated using R-VGA-Whittle without blocking (denoted as “None”) and with blocking starting at different cutoff points for several block sizes (coloured lines), HMC-Whittle (dashed lines), and HMC-exact (dotted lines). True parameter values are denoted using vertical dashed lines.
@@
Figure S5: Posterior densities of the parameters of the univariate SV model estimated using R-VGA-Whittle without blocking (denoted as “None”) and with blocking starting at different cutoff points for several block sizes (coloured lines), HMC-Whittle (dashed lines), and HMC-exact (dotted lines). True parameter values are denoted using vertical dashed lines.
@@
In this section, we study the robustness of R-VGA-Whittle posterior distributions to different values of $S$, the Monte Carlo sample size used to approximate the expectations of the gradient and Hessian in Algorithm 2. We consider sample sizes $S=100,S=500$ and $S=1000$. We apply blocking as in Algorithm 3, and fix the block length to $B=100$. We set $\tilde{n}=176$ for the linear Gaussian SSM, and $\tilde{n}=75$ for the univariate SV model, which corresponds to the 3dB cutoff point as discussed in Section S3.
Figures 6(a)6(c) show the posterior densities of the parameters of the linear Gaussian SSM from 10 runs of R-VGA-Whittle with $S=100$, $S=500$, and $S=1000$, respectively, while Figures 7(a)7(c) show the corresponding plots for the SV model. As the Monte Carlo sample size increases, the variance in the posterior densities across different runs decreases. When $S=1000$, the variance across different runs is significantly reduced, so we choose this as our Monte Carlo sample size for all examples in Section 3 of the main paper.
@@
Figure S6: Posterior densities of the parameters of the linear Gaussian SSM from 10 runs of R-VGA-Whittle (coloured solid lines) with different Monte Carlo sample sizes $S$, from a single run of HMC-Whittle (dashed lines), and from a single run of HMC-exact (dotted lines). True parameter values are marked with dashed lines.
@@
Figure S7: Posterior densities of the parameters of the univariate SV model from 10 runs of R-VGA-Whittle (coloured solid lines) with different Monte Carlo sample sizes $S$, from a single run of HMC-Whittle (dashed lines), and from a single run of HMC-exact (dotted lines). True parameter values are marked with dashed lines.
@@
This section shows the HMC-exact and HMC-Whittle trace plots and effective sample size (ESS) for the parameters in all the models we consider. Figure S8 shows trace plots for the linear Gaussian SSM, Figures S9 and S10 show trace plots for the univariate and bivariate SV models with simulated data, and Figures S11 and S12 show trace plots for the univariate and bivariate SV models with real data.
In general, we find that HMC-Whittle tends to produce posterior samples with much lower correlation than HMC-exact. For the same number of posterior samples, HMC-Whittle has higher ESS than HMC-exact, and the ESS of HMC-Whittle is high for all paramaters, while HMC-exact tends to produce higher ESS for the autoregressive parameters ($\phi$ for univariate examples and $\mathbf{\Phi}$ for bivariate examples) than for the variance parameters ($\sigma_{\eta}$ and $\sigma_{\epsilon}$ for the linear Gaussian model, $\sigma_{\eta}$ for the univariate SV model, and $\mathbf{\Sigma}_{\eta}$ for the bivariate SV model). This is because HMC-Whittle only targets the posterior density of the parameters, whereas HMC-exact targets the joint posterior density of the parameters and states, which are high dimensional. See Table S1 for the ESS corresponding to each parameter in our examples.
Example Parameter ESS (HMC-exact) ESS (HMC-Whittle)
Linear Gaussian $\phi$ 2452.8033 8884.775
Linear Gaussian $\sigma_{\eta}$ 846.3039 7933.656
Linear Gaussian $\sigma_{\epsilon}$ 697.3547 8163.235
Univariate SV (sim. data) $\phi$ 230.00751 6854.622
Univariate SV (sim. data) $\sigma_{\eta}$ 87.94789 6949.066
Bivariate SV (sim. data) $\Phi_{11}$ 814.3441 13064.54
Bivariate SV (sim. data) $\Phi_{22}$ 282.3187 12967.59
Bivariate SV (sim. data) $\Sigma_{\eta_{11}}$ 346.6532 13382.68
Bivariate SV (sim. data) $\Sigma_{\eta_{21}}$ 440.2904 16141.23
Bivariate SV (sim. data) $\Sigma_{\eta_{22}}$ 138.6995 12034.28
Univariate SV (real data) $\phi$ 304.6549 5920.983
Univariate SV (real data) $\sigma_{\eta}$ 118.8147 5980.517
Bivariate SV (real data) $\Phi_{11}$ 730.7380 9058.774
Bivariate SV (real data) $\Phi_{22}$ 796.0991 9242.738
Bivariate SV (real data) $\Sigma_{\eta_{11}}$ 206.4234 11170.016
Bivariate SV (real data) $\Sigma_{\eta_{21}}$ 161.9922 9989.325
Bivariate SV (real data) $\Sigma_{\eta_{22}}$ 245.5852 11127.356
Table S1: The ESS obtained for all parameters considered in the examples described in the main text. The ESS is shown for both HMC-exact and HMC-Whittle from 20000 posterior samples.
@@
Figure S8: Trace plots for the linear Gaussian SSM with simulated data. Red dashed lines denote the true parameter values.
@@
(b) HMC-exact trace plots (samples thinned by a factor of 100).
(c) HMC-Whittle trace plots (samples not thinned).
Figure S9: Trace plots for the univariate SV model with simulated data. Red dashed lines denote the true parameter values.
@@
Figure S10: Trace plots for the bivariate SV model with simulated data. Red dashed lines denote the true parameter values.
(a) HMC-exact trace plots (samples not thinned).
(b) HMC-exact trace plots (samples thinned by a factor of 100).
(c) HMC-Whittle trace plots (samples not thinned).
Figure S11: Trace plots for the univariate SV model applied to exchange rate data.
@@
Figure S12: Trace plots for the bivariate SV model applied to exchange rate data.

### v3

@@ 3.5 Univariate and bivariate SV models with real data
We apply R-VGA-Whittle to estimate parameters in univariate and bivariate SV models from real data. We first apply it to daily JPY-EUR exchange rate data (Kastner et al.,, 2017) from April 1st, 2005 to August 6th, 2015, which consists of $2650$ observations. Typically, the data used in SV models are de-meaned log-returns, which are computed from the raw exchange rate values as:
@@
where $r_{t}$ is the exchange rate on day $t$, and $y_{t}^{*}$ is the log-return on day $t$. We fit the model (24)–(27) to this time series of de-meaned log-returns with $T=2649$, and use the initial variational distribution specified in (33). For the univariate SV model, we also run MCMC-stochvol with the same settings as in Section 3.2.
The HMC-exact trace plots in Figures 12(a) of the online supplement show that posterior samples of $\phi$ are relatively well-mixed, but that the posterior samples of $\sigma_{\eta}$ are highly correlated. We find it necessary to thin these samples by a factor of 50 to reduce autocorrelation and obtain approximately uncorrelated draws. Similarly, MCMC-stochvol also generates highly autocorrelated samples and requires thinning by a factor of 50. In contrast, the HMC-Whittle approach produces well-mixed posterior samples with low autocorrelation.
We compare the posterior densities from R-VGA-Whittle, HMC-Whittle and HMC-exact in Figure 11. The posterior densities obtained from R-VGA-Whittle are similar, but slightly wider than those obtained from the other three methods.
Figure 11: Posterior distributions from HMC-exact (blue), MCMC-stochvol (black), and approximate posterior distributions from R-VGA-Whittle (red) and HMC-Whittle (yellow) for the univariate SV model fitted using JPY-EUR exchange rate data with $T=2649$. Diagonal panels: Marginal posterior distributions of the parameters. Off-diagonal panels: Bivariate posterior distributions of the parameters.
@@
Figures 13(a)13(c) in Section S5 of the online supplement show the HMC-exact and HMC-Whittle trace plots for the bivariate SV model with exchange rate data. Similar to the simulation study, we find that HMC-exact samples are highly correlated, with samples of parameters in $\mathbf{\Phi}$ less correlated than samples of parameters in $\mathbf{\Sigma}_{\eta}$. We thin the HMC-exact samples by a factor of 50 so that the remaining samples can be treated as approximately uncorrelated. Posterior samples from HMC-Whittle have low correlation and are mixed well for all parameters, so no thinning is necessary.
@@
Table 1 shows that in the univariate case, R-VGA-Whittle is slightly slower than HMC-Whittle, but about twice as fast as MCMC-stochvol, and more than 180 times faster than HMC-exact. In the bivariate case, R-VGA-Whittle is more than 23 times faster than HMC-Whittle and more than 750 times faster than HMC-exact.
Figure 12: Posterior distributions from HMC-exact (blue), and approximate posterior distributions from R-VGA-Whittle (red) and HMC-Whittle (yellow) for the bivariate SV model fitted using GBP-EUR and USD-EUR exchange rate data with $T=2649$. Diagonal panels: Marginal posterior distributions of the parameters. Off-diagonal panels: Bivariate posterior distributions of the parameters.
@@ 4 Conclusion
We develop a novel sequential variational Bayes algorithm that uses the Whittle likelihood, called R-VGA-Whittle, for parameter inference with linear non-Gaussian state space models. The Whittle likelihood is a frequency-domain approximation of the exact likelihood. It is available in closed form and its log can be expressed as a sum of individual log-likelihood frequency components. At each iteration, R-VGA-Whittle sequentially updates the variational parameters by processing one frequency or a block of frequencies at a time. We also demonstrate the use of the Whittle likelihood in HMC (HMC-Whittle), and compare the parameter estimates from R-VGA-Whittle and HMC-Whittle to those of HMC with the exact likelihood (HMC-exact). We show that the posterior densities from the three methods are generally similar, though posterior densities from R-VGA-Whittle and HMC-Whittle tend to be wider compared to those from HMC-exact. As HMC-Whittle only targets the parameters and not the states, it tends to produce posterior samples of the parameters that are less correlated than those from HMC-exact. We also show that R-VGA-Whittle and HMC-Whittle are much more computationally efficient than HMC-exact. R-VGA-Whittle is particularly efficient when dealing with bivariate time series, where in our study it is more than 25 times faster than HMC-Whittle and 750 times faster than HMC-exact. R-VGA-Whittle is also shown to be highly computationally efficient in an example involving a long memory (ARFIMA) latent process with a large sample size, which is challenging to fit using classic likelihood-based approaches.
[…]
theoreticalThere are several avenues for future research. First, the work in this paper is largely computational and empirical, and an important next step is to derive theoretical properties of the R-VGA-Whittle algorithm. Second, R-VGA-Whittle assumes a Gaussian variational distribution; future work could attempt to extend this framework to more flexible variational distributions for approximating skewed or multimodal posterior distributions. Third, some tuning parameters of R-VGA-Whittle (such as the number of damping observations and number of damping steps) are currently user-defined, so an adaptive tuning scheme could be developed to select these parameters. Fourth, as the Whittle likelihood is known to produce biased parameter estimates for finite sample sizes, using R-VGA-Whittle with the debiased Whittle likelihood of Sykulski et al., (2019) may improve parameter estimates.
@@ Acknowledgements
This work was supported by ARC SRIEAS Grant SR200100005 Securing Antarctica’s Environmental Future.
@@
Recall that the periodogram is an estimate of the spectral density and can be viewed as a measure of the relative contributions of the frequencies to the total variance of the process (Percival and Walden,, 1993). For example, Figure S1 shows the periodograms for the linear Gaussian SSM in Section 3.1 and the univariate SV model in Section 3.2. This figure shows that for the examples considered, the power generally decreases as frequency increases. In Figure  S2, the trajectory of $\phi$ moves towards the true parameter value within the first few frequencies, with only minor adjustments afterwards. The trajectories for $\sigma_{\eta}$ and $\sigma_{\epsilon}$ behave in a similar manner, but with more gradual changes. This behaviour can also be observed in Figure S3, in which the trajectories of both parameters in the univariate SV model exhibit large changes in the first few iterations, then stabilise afterwards. Higher frequencies that have little effect on the variational mean can thus be processed together in “blocks”. The R-VGA-Whittle updates can then be made based on the Whittle log-likelihood for a “block” of frequencies, which we compute as the sum of the individual log-likelihood contribution evaluated at each frequency within that block.
@@
Figure S1: Plot of raw periodograms of the data (black) and smoothed periodograms (light red) for the univariate linear Gaussian state space model in Section 3.1 and univariate SV model (with $T=2000$, $\phi=0.99$, and $\sigma_{\eta}=0.4$) in Section 3.2. Vertical lines show the frequency cutoffs at half-power (3dB, red dashed line), one-fifth-power (7dB, blue dotted line) and one-tenth-power (10dB, purple dash-dotted line).
Figure S2: Trajectories of the variational mean for the parameters in the linear Gaussian SSM of Section 3.1, estimated using damped R-VGA-Whittle (Algorithm 2) with no blocking. The true parameter values are marked with dashed lines.
Figure S3: Trajectories of the variational mean for the parameters in the univariate SV model in Section 3.2 (with $T=2000$, $\phi=0.99$, and $\sigma_{\eta}=0.4$), estimated using damped R-VGA-Whittle (Algorithm 2) with no blocking. True parameter values are marked with dashed lines.
It is important to choose an appropriate point to start blocking, as well as an appropriate length for each block of frequencies. To determine the frequency at which to begin blocking, we use the 3dB frequency cutoff, which is the frequency on the power spectrum at which the power drops to approximately half the maximum power of the spectrum. The value 3 comes from the fact that, on the decibel scale, a drop from a power level $P_{1}$ to another level $P_{2}=\frac{1}{2}P_{1}$ results in a drop of 3 decibels. The change in decibels is calculated as
@@
We employ a similar idea here, where we define the “cutoff” frequency as the frequency at which the power drops to half of the maximum power in a smoothed version of the periodogram. Specifically, we seek the frequency $\omega^{*}$ such that $\tilde{I}(\omega^{*})=1/2\max(\tilde{I}(\omega))$, where $\tilde{I}(\omega)$ is the periodogram smoothed using Welch’s method (Welch,, 1967). Briefly, Welch’s method works by dividing the frequencies into equal segments, computing the periodogram for each segment, and then averaging over these segments to obtain a smoother periodogram with lower variance. We also experiment with other cutoffs, such as the frequencies at which the power is one-fifth and one-tenth of the maximum power in the smoothed periodogram, which correspond to a drop of approximately 7dB and 10dB, respectively. We mark these cutoff frequencies in Figure S1.
@@
Figure S4: Posterior densities of the parameters of the linear Gaussian SSM in Section 3.1 estimated using R-VGA-Whittle without blocking (denoted as “None”) and with blocking starting at different cutoff points for several block sizes (coloured lines), HMC-Whittle (dashed lines), and HMC-exact (dotted lines). True parameter values are denoted using vertical dashed lines.
@@
Figure S5: Posterior densities of the parameters of the univariate SV model with $T=2000$ in Section 3.2 estimated using R-VGA-Whittle without blocking (denoted as “None”) and with blocking starting at different cutoff points for several block sizes (coloured lines), HMC-Whittle (dashed lines), and HMC-exact (dotted lines). True parameter values ($\phi=0.99$ and $\sigma_{\eta}=0.4$) are denoted using vertical dashed lines.
@@
In this section, we study the robustness of R-VGA-Whittle posterior distributions to different values of $S$, the Monte Carlo sample size used to approximate the expectations of the gradient and Hessian in Algorithm 2. We consider sample sizes $S=100,S=500$ and $S=1000$. We apply blocking as in Algorithm 3, and fix the block length to $B=100$. We set $\tilde{n}=176$ for the linear Gaussian SSM, and $\tilde{n}=10$ for the univariate SV model, which corresponds to the 3dB cutoff point as discussed in Section S3.
Figures 6(a)6(c) show the posterior densities of the parameters of the linear Gaussian SSM from 10 runs of R-VGA-Whittle with $S=100$, $S=500$, and $S=1000$, respectively, while Figures 7(a)7(c) show the corresponding plots for the univariate SV model with $\phi=0.99$ and $\sigma_{\eta}=0.4$. As the Monte Carlo sample size increases, the variance in the posterior densities across different runs decreases. When $S=1000$, the variance across different runs is significantly reduced, so we choose this as our Monte Carlo sample size for all examples in Section 3 of the main paper.
@@
Figure S6: Posterior densities of the parameters of the linear Gaussian SSM in Section 3.1 from 10 runs of R-VGA-Whittle (coloured solid lines) with different Monte Carlo sample sizes $S$, from a single run of HMC-Whittle (dashed lines), and from a single run of HMC-exact (dotted lines). True parameter values are marked with dashed lines.
@@
Figure S7: Posterior densities of the parameters of the univariate SV model with $T=2000$ in Section 3.2 from 10 runs of R-VGA-Whittle (coloured solid lines) with different Monte Carlo sample sizes $S$, from a single run of HMC-Whittle (dashed lines), and from a single run of HMC-exact (dotted lines). True parameter values ($\phi=0.99$, $\sigma_{\eta}=0.4$) are marked with dashed lines.
@@
This section contains trace plots, effective samples sizes (ESS), and $\hat{R}$ values of HMC-exact and HMC-Whittle for all the models we consider (except the ARFIMA model with Student’s $t$ measurement noise, for which we only run HMC-Whittle). Diagnostics for MCMC-stochvol for the univariate SV example in Section 3.2 of the main paper are also included. The ESS is a measure of the number of uncorrelated posterior samples, and the higher the ESS the better; see Gelman et al., (2013) for further details on how Stan calculates ESS. The $\hat{R}$ statistic proposed by Gelman and Rubin, (1992) measures the ratio of the average variance of samples within each chain to the variance of the pooled samples across chains; an $\hat{R}$ close to 1 indicates that the chains have converged.
Figure S8 shows trace plots for the linear Gaussian SSM in Section 3.1; Figures S9 and S10 show trace plots for the univariate and bivariate SV models with simulated data in Sections 3.2 and 3.3; Figure S11 shows the trace plots from the HMC-Whittle sampler for the state space model with Student’s $t$ errors and latent states following an ARFIMA process in Section 3.4; and Figures S12 and S13 show trace plots for the univariate and bivariate SV models with real data in Section 3.5. Table S1 contains the ESS and $\hat{R}$ corresponding to each parameter in our examples. Table S2 shows the total number of posterior samples after burn-in, for HMC-exact, HMC-Whittle, and MCMC-stochvol.
For all examples, the $\hat{R}$ values for all parameters are very close to 1, indicating good convergence of the MCMC chains. HMC-Whittle consistently produces posterior samples with effective sample sizes (ESS) exceeding 1000 across all examples. In contrast, HMC-exact yields ESS values below 1000 for some of the parameters of the univariate SV model with real data and for the bivariate SV models with both simulated and real data, even when we run with a maximum of $30000$ iterations (two chains). Similarly, MCMC-stochvol produces ESS values below 1000 for some parameters in the univariate SV model, for both simulated and real datasets. HMC-Whittle is more computationally efficient than both HMC-exact and MCMC-stochvol because it targets only the posterior distribution of the parameters, whereas HMC-exact targets the full joint posterior of both parameters and latent states, which is high-dimensional.
Table S1: The effective sample sizes (ESS) and $\hat{R}$ statistics from HMC-exact and HMC-Whittle are reported for all parameters in the examples described in the main text. For the univariate SV examples, results from MCMC-stochvol are also included. Examples for which a method was not run (and thus ESS and $\hat{R}$ are not available) are indicated with a dash (–).
Example Parameter HMC-Whittle HMC-exact MCMC-stochvol
$\hat{R}$ ESS $\hat{R}$ ESS $\hat{R}$ ESS
Linear Gaussian $\phi$ 1.0000 2172 1.0001 4160 - -
$\sigma_{\eta}$ 0.9999 1655 1.0000 1801 - -
$\sigma_{\epsilon}$ 0.9998 1659 1.0004 1350 - -
Univariate SV (sim. data with $\phi$ 1.0006 1709 1.0000 8137 1.0029 1413
$\phi=0.99$ and $\sigma_{\eta}=0.4$) $\sigma_{\eta}$ 1.0002 1710 1.0004 1211 1.0001 945
Bivariate SV (sim. data) $\Phi_{11}$ 1.0006 2490 1.0009 1528 - -
$\Phi_{22}$ 1.0007 2267 1.0001 467 - -
$\Sigma_{\eta_{11}}$ 1.0013 2871 1.0019 568 - -
$\Sigma_{\eta_{21}}$ 1.0007 2886 1.0005 599 - -
$\Sigma_{\eta_{22}}$ 1.0002 2048 1.0001 274 - -
ARFIMA (sim. data) $\phi$ 1.0006 6234 - - - -
$\theta$ 1.0014 2299 - - - -
$d$ 1.0008 6053 - - - -
$\sigma_{\eta}$ 1.0014 2083 - - - -
$\nu$ 1.0016 1637 - - - -
Univariate SV (real data) $\phi$ 1.0049 1166 1.0023 482 1.0003 573
$\sigma_{\eta}$ 1.0019 1150 1.0059 184 1.0007 296
Bivariate SV (real data) $\Phi_{11}$ 0.9999 1665 1.0007 1018 - -
$\Phi_{22}$ 0.9999 1558 1.0004 1233 - -
$\Sigma_{\eta_{11}}$ 1.0001 1882 1.0037 277 - -
$\Sigma_{\eta_{21}}$ 1.0008 1708 1.0050 218 - -
$\Sigma_{\eta_{22}}$ 1.0013 1622 1.0024 295 - -
Table S2: Total number of posterior samples used after burn-in (combined across two chains). For all examples and all methods, a burn-in period of 1000 samples is used.
Example HMC-Whittle HMC-exact MCMC-stochvol
Linear Gaussian 4000 20000 -
Univariate SV (sim. data with $\phi=0.99$ and $\sigma_{\eta}=0.4$) 4000 28000 28000
Bivariate SV (sim. data) 4000 28000 -
ARFIMA (sim. data) 20000 - -
Univariate SV (real data) 4000 28000 28000
Bivariate SV (real data) 4000 28000 -
@@
Figure S8: Trace plots for the linear Gaussian SSM with simulated data in Section 3.1 of the main paper. Red dashed lines denote the true parameter values.
@@
(b) HMC-Whittle trace plots (samples not thinned).
(c) MCMC-stochvol trace plots (samples not thinned).
Figure S9: Trace plots for the univariate SV model with simulated data in Section 3.2 of the main paper. Red dashed lines denote the true parameter values ($\phi=0.99$ and $\sigma_{\eta}=0.4$).
@@
Figure S10: Trace plots for the bivariate SV model with simulated data in Section 3.3 of the main paper. Red dashed lines denote the true parameter values.
(a) HMC-Whittle trace plots (samples not thinned).
(b) HMC-Whittle trace plots (samples thinned by a factor of 10).
Figure S11: Trace plots for the state space model with Student’s $t$ errors and latent states following an ARFIMA process in Section 3.4 of the main paper. Red dashed lines denote the true parameter values.
@@
(d) MCMC-stochvol trace plots (samples thinned by a factor of 50).
Figure S12: Trace plots for the univariate SV model applied to exchange rate data in Section 3.5 of the main paper.
(a) HMC-exact trace plots (samples not thinned).
(b) HMC-exact trace plots (samples thinned by a factor of 50).
(c) HMC-Whittle trace plots (samples not thinned).
Figure S13: Trace plots for the bivariate SV model applied to exchange rate data in Section 3.5 of the main paper.
S6 Additional results for the univariate SV model in Section 3.2
This section provides additional figures for an experiment in which we simulate 100 datasets, each with length $T=2000$, from the univariate SV model in Section 3.2 for each value of $\phi\in\{0.7,0.8,0.9,0.99\}$ and $\sigma_{\eta}=0.2$. The number of posterior draws for HMC-Whittle and HMC-exact are kept the same as in Section 3.2 (see Table S2 in this online supplement for a summary). For each parameter setting, the $\hat{R}$ and ESS from HMC-Whittle and HMC-exact are reported in Table S3. In general, the ESS for HMC-Whittle are consistently above 1000 for all parameter settings, while those from HMC-exact tend to be much lower, especially when $\phi=0.7$ and $\phi=0.8$.
Figure 7 in the main paper and Figures S14S16 show comparisons of the 95% credible intervals from R-VGA-Whittle and HMC-exact (left) and R-VGA-Whittle and HMC-Whittle (right) for the first 25 simulated datasets for varying $\phi$. These figures show that the widths of the 95% credible intervals obtained from R-VGA-Whittle and HMC-Whittle are similar, whereas those from HMC-exact are slightly different.
(a) HMC-exact and R-VGA-Whittle.
(b) HMC-Whittle and R-VGA-Whittle.
Figure S14: Comparison of the 95% credible intervals from R-VGA-Whittle and HMC-exact (left) and R-VGA-Whittle and HMC-Whittle (right) for the first 25 simulated datasets from the univariate SV model with $T=2000$, $\phi=0.7$, and $\sigma_{\eta}=0.2$. Posterior means are marked with points.
(a) HMC-exact and R-VGA-Whittle.
(b) HMC-Whittle and R-VGA-Whittle.
Figure S15: Comparison of the 95% credible intervals from R-VGA-Whittle and HMC-exact (left) and R-VGA-Whittle and HMC-Whittle (right) for the first 25 simulated datasets from the univariate SV model with $T=2000$, $\phi=0.8$ and $\sigma_{\eta}=0.2$. Posterior means are marked with points.
(a) HMC-exact and R-VGA-Whittle.
(b) HMC-Whittle and R-VGA-Whittle.
Figure S16: Comparison of the 95% credible intervals from R-VGA-Whittle and HMC-exact (left) and R-VGA-Whittle and HMC-Whittle (right) for the first 25 simulated datasets from the univariate SV model with $T=2000$, $\phi=0.99$ and $\sigma_{\eta}=0.2$. Posterior means are marked with points.
Table S3: $\hat{R}$ and ESS for HMC-Whittle and HMC-exact, averaged over 100 simulations of length $T=2000$, where the true values of $\phi$ are in the set $\{0.7,0.8,0.9,0.99\}$, and $\sigma_{\eta}=0.2$.
True $\phi$ HMC-Whittle HMC-exact
$\hat{R}$ ESS $\hat{R}$ ESS
$\phi$ $\sigma_{\eta}$ $\phi$ $\sigma_{\eta}$ $\phi$ $\sigma_{\eta}$ $\phi$ $\sigma_{\eta}$
0.7 1.0009 1.0008 1328 1525 1.0200 1.0294 130 64
0.8 1.0009 1.0008 1271 1475 1.0149 1.0226 134 67
0.9 1.0007 1.0007 3263 3584 1.0064 1.0100 199 116
0.99 1.0009 1.0006 1407 1405 1.0007 1.0028 447 447